How does 25/125 turn into .20
Answer by Alan3354(69443)

How does 25/125 turn into .20
--------
Any of several ways.
You can do "long division," use a calculator, or
divide NUM and DEN by 25 --> 1/5
(1/5)*(2/2) = 2/10 = 0.2